All ISS systems continue to function nominally, except those noted previously or below. Sunday – off-duty day for CDR Padalka, FE-1 Barratt & FE-2 Wakata. Ahead: Week 6 of Increment 19.
Michael Barratt & Koichi Wakata began the last day of their third session with the NASA/JSC experiment NUTRITION w/Repository. For both, 24-hr. urine collections ended this morning, and Mike finished his session with his blood draw from an arm vein, assisted in the phlebotomy by the FE-2. For Wakata, this NUTRITION/Repository activity was his FD60 session. [The NUTRITION project is the most comprehensive in-flight study done by NASA to date of human physiologic changes during long-duration space flight. It includes measures of bone metabolism, oxidative damage, nutritional assessments, and hormonal changes, expanding the previous Clinical Nutritional Assessment profile (MR016L) testing in three ways: Addition of in-flight blood & urine collection (made possible by supercold MELFI dewars), normative markers of nutritional assessment, and a return session plus 30-day (R+30) session to allow evaluation of post-flight nutrition and implications for rehabilitation.]
Padalka performed the periodic maintenance of the active Russian BMP (Harmful Impurities Removal System) by starting the "bake-out" cycle to vacuum on absorbent bed #1 of the regenerable dual-channel filtration system. The process will be terminated tonight at ~5:15pm EDT before crew sleep, followed tomorrow by Bed #2 regeneration. (Last time done: 4/21-4/22). [Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. The BMP’s regeneration cycle is normally done every 20 days.]
Afterwards, the CDR completed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM (Service Module) and FGB (Funktsionalnyi-Grusovoi Blok), including the weekly collection of the toilet flush (SP) counter and water supply (SVO) readings for calldown to TsUP-Moscow. Additionally, the CDR checked up on the Russian POTOK-150MK (150 micron) air filter unit of the SM’s SOGS air revitalization subsystem, gathering weekly data on total operating time & “On” durations for reporting to TsUP-Moscow. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]
Gennady also conducted the periodic checkup behind panel 139 in the SM on a fluid connector (MNR-NS) of the SM-U urine collection system, looking for potential moisture.
Mike performed the daily procedure of flushing the PWD (Potable Water Dispenser) ambient line with ~50mL of water (into a towel/Ziploc bag). PWD water is currently cleared only for hygienic use, pending Russian approval.
The FE-1 also conducted the daily status check on the BCAT (Binary Colloidal Alloy Test) science payload, running by itself since 5/2. [The status check, conducted on the last image taken by the DCS 760 digital still camera which is controlled by EarthKAM software on an A31p laptop, is to verify proper image focus and camera alignment. The SSC (Station Support Computer) is taking photography of the phase separation occurring in the BCAT Sample 3, with the photo flash going off every half hour.]
In the US Lab, the FE-2 worked on the ITCS (Internal Thermal Control System), mating two MTL (Moderate Temperature Loop) umbilicals (for Supply & Return) to the CHeCS (Crew Health Care System) rack at its UIP (Utility Interface Panel, loc. LAB1S4) to provide cooling to the rack.
Today’s PFCs (Private Family Conferences) for the crew, via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop), were scheduled for Koichi at ~2:40am & ~11:25am, and Mike at ~4:15pm.
The crew completed their regular daily 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (CDR, FE-1, FE-2), ARED advanced resistive exercise device (FE-1, FE-2) and VELO bike with bungee cord load trainer (CDR).
At ~4:50am, Mike Barratt powered up the SM's amateur radio equipment (Kenwood VHF transceiver with manual frequency selection, headset, & power supply) and at ~4:55am conducted a ham radio session with students at the Flemish Space Days, Leuven, Belgium. [For the 5th time, VRI – Flemish Space Industries – has organized the Flemish Space Days, with the support of the Federal and Flemish government. This year the event focuses especially on the new mission of the Belgian ESA astronaut Frank De Winne and on the International Year of Astronomy. In the Brabanthal in Leuven the Flemish space and aviation industry, research centers and educational institutes will show their latest technological and scientific developments to students and the large public. On Friday, 5/8, the Flemish schools were invited in order to try to stimulate young people to choose for a technical or scientific career. Different booths occupied by space and aviation industry, research centers and other actors in the field of aeronautics and astronomy will lead the visitor into the fascinating world of space. Workshops and lectures will highlight the applications of space technology in daily life.]
Weekly Science Update (Expedition Nineteen -- Week 5)
3-D SPACE: Koichi completed his third session of 3D-SPACE. Thank you for performing this experiment.
AgCam (Agricultural Camera): Planned.
ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.
ALTEA DOSI (NASA/ASI): Standing by.
BCAT-3/4 (Binary Colloidal Alloy Test 3/4): Feedback from Peter Lu (PI): Dr. Mike, you did a fabulous job over the weekend setting up and photographing sample 4. The focus and illumination are fantastic, as is clear from the visible phase separation already ongoing (which is, incidentally, much faster than that observed for sample 2 over the past few weeks). And, you managed to trap a large bubble beneath the stir bar, as per the procedures, so it's not interfering with our measurement. Feedback from Chaikin and Hollingsworth at NYU regarding the BCAT-4 crystal growth samples: We are seeing evidence of crystals in samples 8 and 9. Sample 8 is showing some red speckle and curvature in the middle of the sample cell and again to the left. Sample 9 has a small red and green scintillating pattern in the bottom right hand corner. Very nice! Yet, the November pictures of sample 8 and sample 9 looked more dramatic, which indicates that we may want to try taking the photographs at the November angles, since the effects of crystals are likely to be even more pronounced now. Seeing crystals growing in samples 8 and 9 is very exciting and has profound implications regarding how it is that order arises out of disorder in nature. We would be interested to know if you see something similar in sample 10, which has not been captured by the camera. These three samples contain hard spheres, which are monodisperse, polydisperse, and monodisperse seeded with ‘nanodirt’ (large seed particles).”
Biological Rhythms (JAXA): Ongoing. Two runs are for Barratt, two are planned for De Winne.
BISE (CSA, Bodies in the Space Environment): “Thanks Mike for performing a 4th BISE session as voluntary science. Looking forward to see you again in future sessions toward the end of your flight.”
CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Ongoing.
CFE (Capillary Flow Experiment): Reserve.
CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Complete.
CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.
CIR (Combustion Integrated Rack): Nothing new.
CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.
Commercial 2 (JAXA): Cosmo Flower mission was completed by Wakata (slipped from I-18)
Commercial 3 (JAXA): Completed.
CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Complete.
DomeGene (JAXA): Complete.
EarthKAM (Earth Knowledge Acquired by Middle School Students): Planned.
EDR (European Drawer Rack): The rack is continuously active in support of the PCDF (Protein Crystallization Diagnostic Facility) experiment. Two anomalies are currently under investigation.
ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.
ENose (Electronic Nose): Operating.
EPM (European Physiology Module): Underway.
EPO (Educational Payload Operations): Reserve.
EPO J-Astro Report (JAXA): Ongoing.
EPO Space Clothes (JAXA): Complete.
EPO Hiten (Dance, JAXA): Complete.
EPO Moon Score (JAXA): Planned.
EPO Try Zero-G (JAXA): Planned.
EPO Kibo Kids Tour (JAXA): Complete.
EPO Spiral Top (JAXA): Complete.
ETD (Eye Tracking Device): Completed.
EuTEF (European Technology Exposure Facility): Platform nominal. DOSTEL, EXPOSE, FIPEX, On-going science acquisition – nominal. MEDET: Acquired science nominally until 5/6 and then was switched off because the instrument was getting too cold. On-going science acquisition - nominal;
DEBIE-2: the instrument continues to generate empty science packets at regular intervals (of 30 to 34hrs). Science acquisition is pursued with regular power cycling of the instrument (work-around).
FACET (JAXA): “Half of the planned experiments (of 20) are finished,-- with good data.”
FSL (Fluid Science Laboratory): Activated for MMA measurements for 32P undocking. Checkout of FSL optical modes ongoing.
GEOFLOW: No report.
HDTV System (JAXA): To be launched by HTV1.
Holter (JAXA): One of two sessions for FE-2 is slipped from Increment 18. Scheduled on 5/21.
HQPC (JAXA): To be launched by 34P.
ICE CRYSTAL (JAXA): Complete.
IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.
InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): Complete.
Integrated Immune: Complete.
KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.
LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Completed.
Marangoni Experiment for ISS in JAXA FPEF (Fluid Physics Experiment Facility): In progress.
MAXI (JAXA): Exposed Payload, to be launched on 2JA.
Micro-G Clay (JAXA EPO): Complete.
MISSE (Materials ISS Experiment): Ongoing.
Moon Photography from ISS (JAXA EPO): One run performed on the last day of Increment 18.
MSG-SAME (Microgravity Science Glovebox): Complete.
MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.
NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.
NUTRITION w/REPOSITORY: Ongoing.
PADLES (JAXA, Area PADLES 3; Passive Dosimeter for Lifescience Experiment in Space): Continuing radiation dose accumulation.
PCDF-PU (Protein Crystallization Diagnostic Facility - Process Unit): Growth cycle, EP4 Cycle4 rerun ended on 5/5. In this reactor, science team sees about 10 large crystals that will be kept to be downloaded with PCDF-PU. Additionally in this reactor, the science team observed nucleation occurring at a higher temperature than in ground testing. The cause for this is still being studied.
Nucleation cycle EP1 Cycle3 was started on 5/5. This cycle will conclude on 5/8. This cycle studies nucleation at a certain precipitant and protein concentration but at three different temperature rampings.
PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete.
PMDIS (Perceptual Motor Deficits in Space): Complete.
POLCA/GRAVIGEN (ESA): Complete.
RadGene & LOH (JAXA): Complete.
SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.
SEDA-AP (JAXA): Exposed Payload,- to be launched by 2JA.
SHERE (Shear History Extensional Rheology Experiment): Complete.
SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Mike, thanks for completing your first week of sleep logging. You will have another week scheduled in a few weeks. We are targeting every third week. Thanks for the continued support of the experiment.”
SMILES (JAXA): Exposed Payload, to be launched by HTV1.
SOLAR (Solar Monitoring Observatory): Awaiting the start of the next Sun visibility window which is expected to begin on 5/23.
SOLO (Sodium Loading in Microgravity): No report.
SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): Reserve.
SPICE (Smoke Point In Co-flow Experiment): Restarted SPICE operations on GMT 4/29.
Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.
TRAC (Test of Reaction & Adaptation Capabilities): Planned.
VLE (Video Lessons ESA): VLE-1 completed.
WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Complete/Planned (2J/A Stage).
CEO (Crew Earth Observations): “Through 5/4, the ground has received a total of 5,499 of ISS CEO imagery for review and cataloguing. “Many of your imagery are under review and we are now pleased to confirm your acquisition of views of the following targets: Houston, TX; Delhi, India; Paris, France; Lake Poopo, Bolivia; Lake Chad, Chad; Galapagos Islands, Ecuador; and Kilauea and Mauna Loa, Hawai’i. Nice going! Also you are being very responsive to our specific requests with better than 20% of your imagery being acquired at target times provided. Your excellent, well-focused view of Ankara, Turkey will be posted on NASA/GSFC’s Earth Observatory website this weekend. Among the best recent images from the ISS of the Turkish capital city, your imagery beautifully documents this ancient and complex urban area located in an active earthquake region. Kudos to the crew for a fine image of a requested target!”
CEO Note: In recent days, ISS daylight-awake orbit tracks have shifted rapidly into the Southern Hemisphere which is now some six weeks into the fall season, and both day length and sun elevation are significantly lowering. This situation along with deteriorating seasonal weather greatly limits good view opportunities for targets. The situation is further aggravated by the fact that the ISS orbit tracks nearly parallel with the terminator. The consequence is very low light right of track, low light near nadir, and adequate to good light left of track. Beginning today and for the next 5 to 7 days, there may be no targets with suitable illumination or weather.
CEO photography can be studied at this “Gateway” website:
http://eol.jsc.nasa.gov (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).
Look-ahead: Progress M-02M/33P Docking Events (Tuesday, 5/12; times EDT):
· Start Flyaround – 2:51:50pm (at ~400m)
· Start Stationkeep – 3:00:50pm (at ~160m, directly below, i.e. nadir, of DC1)
· Start Final Approach – 3:12pm
· Local Sunset – 3:17:28pm (Progress at 92m)
· Docking at DC1 nadir port – 3:23pm
· Local Sunrise – 3:34pm
· Hooks closed – 3:43pm.
Significant Events Ahead (all dates Eastern Time, some changes possible!):
05/11/09 -- STS-125/Atlantis Hubble Service Mission 4 (SM4) -- fifth service mission to HST (2:01pm EDT)
05/12/09 -- Progress M-02M/33P docking (~3:23pm EDT)
05/18/09 -- Progress M-01M/32P deorbit (~3:00pm EDT)
05/27/09 -- Soyuz TMA-15/19S launch
05/29/09 -- Soyuz TMA-15/19S docking (FGB nadir)
Six-person crew on ISS
06/05/09 -- Russian EVA-22
06/10/09 -- Russian EVA-23
06/13/09 -- STS-127/Endeavour/2J/A launch - JEM EF, ELM-ES, ICC-VLD
07/17/09 – Progress M-02M/33P undock & deorbit
07/20/09 -- Soyuz TMA-14/18S relocation (from SM aft to DC1)
07/24/09 -- Progress 34P launch
07/26/09 -- Progress 34P docking (SM aft)
08/06/09 -- STS-128/Discovery/17A – MPLM (P), LMC
09/01/09 -- H-IIB (JAXA HTV-1) launch – tentative
09/07/09 -- H-IIB (JAXA HTV-1) berth
09/30/09 -- Soyuz TMA-16/20S launch
10/02/09 -- Soyuz TMA-16/20S docking (SM aft, until MRM2 w/new port)
10/08/09 -- H-IIB (JAXA HTV-1) unberth
10/11/09 – Soyuz TMA-14/18S undock
10/15/09 -- Progress 35P launch
11/10/09 -- 5R/MRM2 (Russian Mini Research Module, MIM2) on Proton -- tentative
11/12/09 -- STS-129/Atlantis/ULF3 - ELC1, ELC2
12/07/09 -- Soyuz TMA-17/21S launch
12/26/09 -- Progress 36P launch
02/03/10 -- Progress 37P launch
02/??/10 -- STS-130/Endeavour/20A – Node-3 + Cupola -- tentative
02/11/10 -- STS-131/Atlantis/19A – MPLM(P), LMC -- tentative
04/02/10 -- Soyuz TMA-18/22S launch
04/08/10 -- STS-132/Discovery/ULF4 – ICC-VLD, MRM1 -- tentative
04/27/10 -- Progress 38P launch
05/29/10 -- Soyuz TMA-19/23S launch
05/31/10 -- STS-133/Endeavour/ULF5 – ELC3, ELC4 -- tentative
06/??/10 – ATV2 – Ariane 5 (ESA)
06/25/10 -- Progress 39P launch
08/11/10 -- Progress 40P launch
09/29/10 -- Soyuz TMA-20/24S launch
12/??/11 -- Proton 3R/MLM w/ERA.
10/19/10 -- Progress 41P launch
12/??/11 – 3R Multipurpose Laboratory Module (MLM) w/ERA – on Proton